# AdaSCALE: Adaptive Scaling for OOD Detection

### v1

@@ 4.2 Revisiting Static Scaling Mechanisms
4.3 Proposed Approach: Adaptive Scaling
Building on our observations, we propose AdaSCALE (Adaptive SCALE), a novel approach that introduces dynamic, sample-specific adjustments to the scaling procedure. The key insight of AdaSCALE is that the percentile threshold should be a function of each test sample’s estimated OOD likelihood rather than a fixed value. The scaling factor increases as the percentile threshold rises (i.e., when more activations are excluded from the denominator in Equation 10). For optimal ID-OOD separation, we must scale ID samples more strongly than OOD samples, requiring a higher percentile for ID samples. Leveraging as an OOD likelihood estimator, we can now define an adaptive percentile threshold as:
where and are hyperparameters that define the minimum and maximum limits of percentile threshold. It ensures that samples with lower OOD likelihood (i.e., likely ID samples) receive higher percentile thresholds, resulting in stronger scaling.
Algorithm 1 summarizes our adaptive scaling factor computation procedure. We implement two variants: AdaSCALE-A scales activations as , similar to ASH [7] and SCALE [54]. AdaSCALE-L scales logits as , similar to LTS [8]. For OOD detection, we use the energy score on (directly or indirectly) scaled logits, with higher values indicating higher ID likelihood. This approach enables per-sample dynamic scaling, as outlined in Figure 2.
5 Experiments
We use pre-trained models provided by PyTorch for ImageNet-1k experiments. For CIFAR experiments, we train three models per network using the standard cross-entropy loss and report the mean results across these three independent trials. The evaluation setup is provided in Table 2.
Conventional OOD detection ID datasets Near-OOD Far-OOD Network CIFAR-10/100 CIFAR-100/10 MNIST, SVHN, WRN-28-10, DenseNet-101 TIN Textures, Places365 ImageNet-1k SSB-Hard iNaturalist, Places, EfficientNetV2-L, ResNet-101 NINCO OpenImage-O DenseNet-201, ViT-B-16 ImageNet-O Textures ResNet-50, ResNeXt-50 Places RegNet-Y-16, Swin-B Covariate shifted datasets for full spectrum OOD detection ImageNet-1k ImageNet-C, ImageNet-R, ImageNet-V2, ImageNet-ES
Baselines. We consider the following post-hoc methods: MSP [16], ODIN [30], EBO [33], ReAct [47], MLS [20], ASH [7], SCALE [54], BFAct [25], LTS [8], OptFS [64]. Currently, SCALE is the best-performing method (with ResNet-50), while OptFS is the best-generalizing method.
Hyperparameters. The hyperparameters are determined via automatic parameter search provided by OpenOOD [57, 61]. The best results are bold, and the second-best results are underlined across all results. We refer readers to Appendices A, B, C D, and E for complete details and studies.
5.1 Empirical Results
ImageNet-1k benchmark:
The ImageNet-1k benchmark is crucial for OOD evaluation, as large-scale datasets offer a comprehensive representation of real-world scenarios. Consequently, post-hoc methods are particularly valuable when training data is inaccessible and retraining is computationally prohibitive. We compare our proposed method, AdaSCALE, with recent state-of-the-art approaches across eight architectures on the ImageNet-1k benchmark, as presented in Table 1. AdaSCALE demonstrates consistently strong performance across all architectures compared to existing methods. Specifically, it surpasses the best-generalizing method, OptFS, by in the FPR@95/AUROC metric for near-OOD detection across all architectures. Additionally, it outperforms the best-performing method, SCALE (on ResNet-50), by in the same metric. A closer observation reveals that while OptFS excels in architectures such as EfficientNet, ViT-B-16, and Swin-B, scaling baselines perform comparably or even better in architectures like ResNet-50, ResNet-101, RegNet-Y-16, and DenseNet-201. In contrast, AdaSCALE-A achieves the best performance in near-OOD detection across all architectures, except for Swin-B, where BFAct performs optimally. Furthermore, the effectiveness of AdaSCALE extends beyond the near-OOD detection to far-OOD detection, demonstrating an average gain of 21.67% in the FPR@95 metric compared to OptFS.
Method Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC MSP 86.75 50.13 74.28 65.67 MLS 88.93 49.48 76.33 65.06 EBO 89.11 49.08 77.21 63.44 ReAct 87.22 51.38 67.23 69.53 ASH 87.01 52.02 72.36 65.70 SCALE 86.75 52.27 69.36 68.97 BFAct 87.12 51.14 66.13 69.69 LTS 86.46 53.29 66.76 71.63 OptFS 85.83 52.17 63.32 71.44 AdaSCALE-A 81.34 55.03 58.87 72.41 AdaSCALE-L 81.62 55.14 59.19 72.85
Full-Spectrum OOD Detection:
Full-spectrum OOD (FS-OOD) detection extends conventional OOD detection by incorporating model’s ability to generalize on covariate-shifted ID inputs. We present FS-OOD detection results in Table 3. We can observe that this is a highly challenging task, as covariate-shifted ID datasets cause a significant performance drop for all methods compared to the conventional case. Despite this, AdaSCALE outperforms OptFS by 4.49 and 4.13 points on average in the FPR@95 metric for full-spectrum OOD detection across both near- and far-OOD datasets.
ISH training:
Apart from enhancing the prior postprocessor ASH [7], SCALE [56] introduces a training regularization to emphasize samples with more distinct ID characteristics. We assess the performance of each method in ResNet-50 and ResNet-101 model following this regularization in Table 4. The results indicate that AdaSCALE maintains a substantial advantage, surpassing the second-best method, SCALE, by and in FPR@95 / AUROC for near- and far-OOD detection in ResNet-50, respectively. Moreover, AdaSCALE demonstrates superior performance beyond conventional OOD detection, with corresponding improvements of 4.10%/5.87% and 9.70%/1.12% in a full-spectrum setting. Furthermore, ISH regularization further amplifies the performance gap between AdaSCALE-A and OptFS, enhancing the near-OOD detection improvement from 12.96% / 6.44% to 15.18%/14.83%. These findings also seem to generalize to the ResNet-101 network, as shown in Table 4.
Method OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD ResNet-50 MSP 74.07 / 62.16 51.13 / 84.64 87.52 / 40.36 74.41 / 61.52 MLS 74.38 / 66.43 41.57 / 88.90 88.89 / 39.49 71.53 / 61.69 EBO 74.68 / 66.46 41.85 / 88.83 89.05 / 39.18 71.77 / 61.11 ReAct 71.98 / 70.81 28.76 / 93.49 87.78 / 43.88 61.87 / 71.64 ASH 67.99 / 73.46 23.88 / 94.67 85.74 / 45.29 57.81 / 72.81 SCALE 65.68 / 76.41 20.77 / 95.62 84.31 / 48.40 54.48 / 74.79 BFAct 71.59 / 70.85 28.38 / 93.50 87.51 / 43.95 61.39 / 71.43 LTS 66.32 / 75.03 22.07 / 95.28 85.08 / 46.16 57.11 / 73.06 OptFS 67.71 / 73.03 24.65 / 94.18 85.38 / 45.91 57.09 / 72.95 AdaSCALE-A 57.43 / 80.86 16.52 / 96.46 80.85 / 51.24 51.57 / 75.63 AdaSCALE-L 56.83 / 80.81 17.62 / 96.22 80.97 / 50.59 53.43 / 74.62 ResNet-101 MSP 71.39 / 68.31 51.00 / 84.81 85.70 / 45.72 73.69 / 62.79 MLS 72.32 / 71.94 41.04 / 88.99 87.39 / 44.88 69.94 / 63.23 EBO 72.78 / 71.92 41.45 / 88.87 87.64 / 44.66 70.23 / 62.65 ReAct 67.74 / 75.74 28.53 / 93.47 85.40 / 48.83 60.50 / 72.18 ASH 66.03 / 77.79 25.21 / 94.43 83.95 / 50.57 56.91 / 73.37 SCALE 64.30 / 78.98 23.09 / 94.95 83.14 / 51.05 55.88 / 73.27 BFAct 67.53 / 75.86 28.32 / 93.40 85.11 / 48.96 60.12 / 71.83 LTS 66.32 / 75.03 22.07 / 95.28 85.08 / 46.16 57.11 / 73.06 OptFS 67.71 / 73.03 24.65 / 94.18 85.38 / 45.91 57.09 / 72.95 AdaSCALE-A 54.66 / 83.52 16.81 / 96.32 78.52 / 55.19 49.92 / 76.20 AdaSCALE-L 53.91 / 83.49 17.55 / 96.15 78.63 / 54.60 51.47 / 75.47
CIFAR benchmark:
CIFAR benchmarks are traditionally used to evaluate OOD detectors in smaller-scale settings. We compare AdaSCALE with recent state-of-the-art post-hoc baselines on CIFAR benchmarks using WRN-28-10 and DenseNet-101 networks, reporting the averaged performance over both architectures in Table 5. AdaSCALE outperforms all methods in average AUROC metric across CIFAR benchmarks in both near- and far-OOD detection. For far-OOD detection on CIFAR-10 benchmark, AdaSCALE-A achieves the best FPR@95 score of 33.11, outperforming the MSP baseline by approximately 1.4 points. Similarly, AdaSCALE-A attains the best FPR@95 / AUROC of 43.07 / 90.31 in near-OOD detection, though MSP remains competitive. In near-OOD detection on CIFAR-100 benchmark, AdaSCALE-A achieves the highest AUROC of 81.35, while in far-OOD detection, AdaSCALE-L reaches the best performance with FPR@95 / AUROC of 52.49 / 82.21. While activation-shaping methods (ReAct, ASH, SCALE, OptFS) perform remarkably well in comparison to logit-based methods in a large-scale setting (ImageNet-1k), they generally seem to underperform in small-scale settings (CIFAR). In contrast, AdaSCALE’s achieves consistently superior performance across all setups, demonstrating its robustness.
@@ 5.1 Empirical Results
Accuracy:
5.2 Ablation / hyperparameter studies
Sensitivity of , , and . As discussed in Sec. 4.1.2, ID and OOD distinctions are more pronounced in high activations but diminish as more activations are included. We analyze the impact of (used in activation shift) and (used in perturbed activation) using a heatmap of validation AUROC with ResNet-50 (Figure 6(a)). The darker region indicates higher AUROC, suggesting optimal values of and for ResNet-50 model. Furthermore, the heatmap suggests that is far more critical hyperparameter than . The hyperparameter controls the weighting of in computing , the predetermined OOD likelihood. The sensitivity analysis is presented in Figure 6(b) which shows near-OOD and far-OOD detection using FPR@95. It suggests optimal FPR@95 is achieved at .
Predetermined OOD likelihood .
@@ 5.2 Ablation / hyperparameter studies
Adaptive scaling depends on predetermined OOD likelihood to determine the extent of scaling. We study the effect of various predetermined OOD likelihood functions on OOD detection using ResNet-50 network (ImageNet-1k) in Table 6. It clearly shows Q component of being most critical while as correction term being a relatively superior choice.
OOD Detection Near-OOD Far-OOD Q 59.43 / 78.14 19.70 / 95.73 70.39 / 74.00 21.40 / 95.31 58.97 / 78.98 17.84 / 96.14 65.11 / 76.23 19.76 / 95.67 58.91 / 78.74 18.02 / 96.08
Adaptive percentile.
Unlike SCALE which uses constant percentile, AdaSCALE uses dynamic percentile lying in range adaptive to each sample. We show the effect of various percentile limit ranges in Figure 7 in the form of a heatmap on the AUROC validation metric on ImageNet-1k benchmark. The extent of darkness in the heatmap conveys a strong performance (corresponding to highest AUROC / lowest FPR@95). The diagonal entries, representing the results of SCALE, are lighter in comparison to the rest of the cells, denoting the results of AdaSCALE. Hence, it can be observed that using adaptive percentile leads to better OOD detection in comparison to static percentile.
ID statistics.
@@
With the rise of large models, where training data is often undisclosed or inaccessible, relying on full training ID datasets for OOD detection has become increasingly impractical. While our approach leverages ID statistics for optimal performance, we rigorously assess its effectiveness with limited data. We conduct experiments on ImageNet-1k using ID samples to compute ID statistics, where . The results in Table 7 confirm that even with substantially restricted access to ID data, AdaSCALE-A achieves state-of-the-art performance.
OOD Detection FS-OOD Detection Near-OOD Far-OOD Near-OOD Far-OOD 10 59.69 / 78.52 18.25 / 96.03 81.66 / 49.76 52.18 / 75.01 100 59.05 / 78.92 17.79 / 96.13 81.52 / 49.19 52.39 / 74.82 1000 58.99 / 78.95 17.86 / 96.13 81.55 / 49.48 52.24 / 74.94 5000 58.97 / 78.98 17.84 / 96.14 81.52 / 49.35 52.33 / 74.89
@@ 6 Conclusion
We propose AdaSCALE, a novel post-hoc OOD detection method that dynamically adjusts the scaling process based on a sample’s estimated OOD likelihood. Leveraging the observation that OOD samples exhibit larger activation shifts under minor perturbations, AdaSCALE assigns stronger scaling to likely ID samples and weaker scaling to likely OOD samples, enhancing ID-OOD separability. AdaSCALE achieves state-of-the-art performance as well as generalization across architectures requiring negligibly few ID samples, making it highly practical for real-world deployment.
Supplementary Material

### v2

@@ 5.1 Empirical Results
CIFAR benchmark. We compare AdaSCALE with post-hoc baselines on CIFAR benchmarks using WRN-28-10 and DenseNet-101 networks, reporting the averaged performance in Table˜3. AdaSCALE outperforms all methods in average AUROC metric across CIFAR benchmarks in both near- and far-OOD detection. For far-OOD detection on CIFAR-10 benchmark, AdaSCALE-A achieves the best FPR@95 score of 33.11, outperforming the MSP baseline by approximately 1.4 points. Similarly, AdaSCALE-A attains the best FPR@95 / AUROC of 43.07 / 90.31 in near-OOD detection, though MSP remains competitive. In near-OOD detection on CIFAR-100 benchmark, AdaSCALE-A achieves the highest AUROC of 81.35, while in far-OOD detection, AdaSCALE-L reaches the best performance with FPR@95 / AUROC of 52.49 / 82.21. While activation-shaping methods perform well in ImageNet-1k, they seem to underperform in CIFAR. In contrast, AdaSCALE achieves consistently superior performance across all setups.
ImageNet-1k benchmark. We compare our proposed method, AdaSCALE, with recent state-of-the-art approaches across eight architectures on the ImageNet-1k benchmark, as presented in Table˜2. AdaSCALE demonstrates consistently strong performance across all architectures compared to existing methods. Specifically, it surpasses the best-generalizing method, OptFS, by in the FPR@95/AUROC metric for near-OOD detection across all architectures. Additionally, it outperforms the best-performing method, SCALE (on ResNet-50), by in the same metric. A closer observation reveals that while OptFS excels in architectures such as EfficientNet, ViT-B-16, and Swin-B, scaling baselines perform comparably or even better in architectures like ResNet-50, ResNet-101, RegNet-Y-16, and DenseNet-201. In contrast, AdaSCALE-A achieves the best performance in near-OOD detection across all architectures, except for Swin-B, where BFAct performs optimally. Furthermore, effectiveness of AdaSCALE extends beyond near-OOD detection to far-OOD detection, demonstrating an average gain of 21.67% over OptFS in the FPR@95 metric.
Method Near-OOD Far-OOD FPR@95 AUROC FPR@95 AUROC ReAct 87.22 51.38 67.23 69.53 ASH 87.01 52.02 72.36 65.70 SCALE 86.75 52.27 69.36 68.97 BFAct 87.12 51.14 66.13 69.69 LTS 86.46 53.29 66.76 71.63 OptFS 85.83 52.17 63.32 71.44 AdaSCALE-A 81.34 55.03 58.87 72.41 AdaSCALE-L 81.62 55.14 59.19 72.85
FSOOD Detection. FSOOD detection extends conventional OOD detection by incorporating model’s ability to generalize on covariate-shifted ID inputs. We present FSOOD detection results in Table˜4. We can observe that this is a highly challenging task, as covariate-shifted ID datasets cause a significant performance drop for all methods compared to conventional case. Despite this, AdaSCALE outperforms OptFS by 4.49 and 4.13 points on average in the FPR@95 metric for FSOOD detection across both near- and far-OOD datasets.
Accuracy: Like SCALE and LTS, AdaSCALE applies linear transformations to scale activations or logits, preserving accuracy, unlike post-hoc rectification methods (ASH, ReAct).
5.2 Ablation / hyperparameter studies / Discussion
Sensitivity study. We study the sensitivity of AdaSCALE-A hyperparameters with ResNet-50 on ImageNet-1k (Table˜7). The optimal perturbation magnitude is , close to RGB means and sufficient to disturb peak activations. Setting ensures activation shift dominates the regularizer . Optimal performance at confirms that shifts in peak activations are the most reliable OOD signal, while empirically performs best. yielding optimal result supports our hypothesis that perturbation should be trivial enough to only disturb peak activations. (See Appendix˜E for complete sensitivity study.)
Method ImageNet-1k CIFAR-100 Near-OOD Far-OOD Near-OOD Far-OOD OptFS 71.91 / 73.88 38.25 / 89.67 76.81 / 73.02 60.23 / 77.76 AdaSCALE-A 62.29 / 79.72 32.72 / 91.82 58.00 / 81.14 56.47 / 80.99
Constrained hyperparameter tuning. The hyperparameters = determined with ResNet-50 model in ImageNet-1k benchmark generalize across architectures and datasets. To demonstrate this, we conduct experiments only allowing to be tuned across 8 and 2 diverse architectures in ImageNet-1k and CIFAR-100 datasets, respectively and present the results in Table˜5. Inspired by early works that tune the percentile hyperparameter in , we set to the lower limit 60. Even under this highly constrained setting, AdaSCALE significantly outperforms the previous state-of-the-art (OptFS) by an average of 13% (FPR95) / 8% (AUROC) on near-OOD and 14% (FPR95) / 2% (AUROC) on far-OOD datasets on ImageNet-1k benchmark. The superiority of AdaSCALE over OptFS is also evident in a small-scale setting (CIFAR-100).
@@ 5.2 Ablation / hyperparameter studies / Discussion
Estimated OODness . Adaptive scaling depends on estimated OODness to determine the extent of scaling. We study the effect of various estimated OODness functions using ResNet-50 network (ImageNet-1k) in Table˜6. It clearly shows Q component of being most critical while as correction term being a relatively superior choice. However, estimated OODness alone – without adaptive scaling – does not result in strong performance.
Near-OOD Far-OOD without scaling 79.81 / 72.32 84.00 / 68.13 59.43 / 78.14 19.70 / 95.73 70.39 / 74.00 21.40 / 95.31 58.97 / 78.98 17.84 / 96.14 65.11 / 76.23 19.76 / 95.67 58.91 / 78.74 18.02 / 96.08
| Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | |||||
| 0.1 | 63.76 / 77.50 | 19.26 / 95.85 | 1 | 67.49 / 75.45 | 20.41 / 95.54 | 1% | 58.97 / 78.98 | 17.84 / 96.14 | 1% | 59.08 / 78.56 | 18.35 / 96.02 | 1% | 61.77 / 78.29 | 19.28 / 95.77 |
| 0.5 | 58.97 / 78.98 | 17.84 / 96.14 | 10 | 58.97 / 78.98 | 17.84 / 96.14 | 10% | 60.40 / 77.89 | 19.72 / 95.63 | 5% | 58.97 / 78.98 | 17.84 / 96.14 | 5% | 58.97 / 78.98 | 17.84 / 96.14 |
| 1.0 | 61.60 / 76.96 | 19.31 / 95.84 | 100 | 59.03 / 78.10 | 19.39 / 95.80 | 100% | 60.99 / 76.49 | 21.88 / 94.96 | 100% | 62.43 / 75.44 | 23.05 / 94.69 | 100% | 67.37 / 73.08 | 22.93 / 95.01 |
| Method | Gaussian blur | JPEG compression | Adversarial training | |||
| Near-OOD | Far-OOD | Near-OOD | Far-OOD | Near-OOD | Far-OOD | |
| SCALE | 72.68 / 70.15 | 29.65 / 92.72 | 69.73 / 71.84 | 27.46 / 93.56 | 66.99 / 76.10 | 24.77 / 94.40 |
| OptFS | 71.84 / 69.66 | 30.27 / 92.24 | 69.43 / 71.12 | 28.34 / 92.95 | 70.11 / 73.51 | 28.32 / 93.42 |
| AdaSCALE-A | 67.53 / 75.18 | 28.11 / 93.31 | 61.24 / 77.59 | 25.87 / 94.06 | 60.21 / 80.68 | 21.13 / 95.35 |
Robustness to real-world image corruptions. To evaluate the practical robustness of our method, we assess its performance when input images are subjected to common real-world corruptions. We compare AdaSCALE against SCALE and OptFS on the ImageNet-1k dataset using a ResNet-50, applying Gaussian blur (kernel=5, sigma=(0.01, 0.5)) and JPEG compression (50%) to the input images. The results (FPR@95/AUROC), presented in Table˜8 without any hyperparameter re-tuning, show that AdaSCALE’s performance advantage is maintained even under these conditions.
Robustness to adversarial training. We evaluate AdaSCALE’s performance on adversarially trained model (=0.05) using the exact same hyperparameters as for the standard ResNet-50. The results (FPR@95 / AUROC) in Table˜8 show that AdaSCALE outperforms baseline methods (SCALE and OptFS). This highlights the robustness of AdaSCALE across diverse training schemes.
Near-OOD Far-OOD 10 59.69 / 78.52 18.25 / 96.03 100 59.05 / 78.92 17.79 / 96.13 1000 58.99 / 78.95 17.86 / 96.13 5000 58.97 / 78.98 17.84 / 96.14
@@
ID statistics. With the rise of large models, where training data is often undisclosed or inaccessible, relying on full training ID datasets for OOD detection has become increasingly impractical. We rigorously assess AdaSCALE’s effectiveness with limited data by conducting experiments on ImageNet-1k with ResNet-50 using ID samples to compute ID statistics, where . The results in Table˜9 confirm that even with substantially restricted access to ID data, AdaSCALE-A achieves state-of-the-art performance.
Fixed percentile (SCALE/LTS) 33 40 45 48 54 Variable percentile (AdaSCALE) 152 149 155 152 164 Latency ratio (AdaSCALE / SCALE) 4.66 3.76 3.42 3.14 3.02
Latency. AdaSCALE incurs extra forward pass to compute perturbed activation . Also, top-k operations (time complexity: ) are applied to and to estimate OODness. Comparing variable vs. fixed percentiles for scaling in Table˜10 over 10,000 trials, we observe that variable percentiles induce higher latency, though the latency ratio decreases with higher-dimensional activation spaces. While SCALE and OptFS have similar latency, AdaSCALE’s is 2.91x higher with gradient attribution for trivial pixel perturbation, dropping to 1.56x with random pixel perturbation.
@@ 6 Conclusion
We propose AdaSCALE, a novel post-hoc OOD detection method that dynamically adjusts the scaling process based on a sample’s estimated OODness. Leveraging the observation that OOD samples exhibit larger activation shifts under minor perturbations, AdaSCALE assigns stronger scaling to likely ID samples and weaker scaling to likely OOD samples, enhancing ID-OOD separability. AdaSCALE achieves state-of-the-art performance as well as generalization across architectures requiring negligibly few ID samples, making it highly practical for real-world deployment. It lays the foundation for future research to build upon the adaptive scaling mechanism.
7 Broader Impacts
This work has positive impact in trustworthy deep learning by enabling detection of OOD samples. Reliable OOD detection can help prevent models from making overconfident predictions in high-stakes applications.
Appendix